OpenStudy (anonymous):

Reba is eating her favorite cereal.  There are 3 2/3 servings remaining in the box.  Reba pours only 1/3 of a serving into her bowl at a time.  How many more bowls can Reba have before the box is empty?

OpenStudy (anonymous):

11??

OpenStudy (texaschic101):

(3 2/3) / (1/3) = (11/3) / (1/3) = 11/3 * 3/1 = 33/3 = 11